OpenStudy (anonymous):

whats 13.33333333 into fraction

OpenStudy (amistre64):

13 3/9

OpenStudy (anonymous):

i need it in a whole

OpenStudy (amistre64):

then convert it ....

OpenStudy (anonymous):

120/9?

OpenStudy (anonymous):

can you tell me how you got 13 3/9 please

OpenStudy (amistre64):

:) maybe .. lets see 13 3/9 = 13 1/3 13(3)+1 39+1 ------- = -------- = 40/3 3 3

OpenStudy (amistre64):

the trick is in the repeating decimal ...

OpenStudy (amistre64):

13 is gonna be 13 regardless so lets look at the .3333... part lets rename it: A = .3333333 ; A*10 = 3.333333 10A = 3.33333 - A = .33333 ---------------- 9A = 3.000000 A = 3/9

OpenStudy (amistre64):

the shortcut is simply take the repeating part; and put that many 9s under it. .12312312312312... ; we see that the 123 part is repeating and it has 3 digits 123 --- is its fraction .. reduce as needed 999

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

13 1/3 or as improper fraction it will be 40/3

OpenStudy (amistre64):

youre welcome :)